The spring under the action of a force of 4 N lengthened by 5 mm. Determine the weight of the weight that will extend this spring by 16 mm.


1
Expert's answer
2021-07-05T15:23:09-0400

Let's first find the spring constant from the Hooke's law:


"F=kx,""k=\dfrac{F}{x}=\dfrac{4\ N}{5\cdot10^{-3}\ m}=800\ \dfrac{N}{m}."

Finally, we can find the weight of the weight that will extend this spring by 16 mm:


"F=kx=800\ \dfrac{N}{m}\cdot16\cdot10^{-3}\ m=12.8\ N."